Had about half a bottle of very clean fog water. Pulled about 40 ml into aliquot. Unfortunately, we are out of sampling bottles, so I unplugged the active sensor. Will clean it once we get ready to sample with new bottles…
We got back up and running Weds 0617 with a new power supply – evening looked promising and fan was on but no sample next morning.
Thurs 0618 installed previous SD card triggering when RH>89% trigger. Drier evening, higher cloud, no trigger.
Today, Fri 0619 we installed the passive sensor, which is currently running (as of 1400 PDT Daniel), though not wired to active sampler yet. Another breakthrough. Have found access to ethernet, will install Tuesday next week.
Forecast for warm, dry weekend, so the sampler is decommissioned until Monday 0622.
Greetings fog-folks, our new active sampling site at SFSU is deployed. This sampler was located at Montara Lighthouse (MTA) last year but moved to SFSU campus to try and find higher fog deposition rates. SFSU is in the SW corner of SF, ~1 mile in from the coast. The sampler is mounted on the roof deck of a 10 story building at ~50 masl. We deployed June 4th, 2015, so here is the summary of operations so far…
060415: Deployed active sampler on rooftop deck, tested system, collected a blank, which Peter took with him. Have not yet mounted passive sensor.
060515-060915: Doors remained closed, suspect droplet sensor is not triggering since RH is often above 75% here. System checked and tested fine. Full lab cleaning.
061015: Rain in night, doors open, fan on but no sample. Rain drops on floor of sampler near door. Assume rain drops were too heavy for fan to draw all the way to strands?? Full lab clean.
061115: Doors open, fan on, sample in bottle ~30-40mL, sample refrigerated (fixed with HCl and refrigerated 061515). Full lab clean.
061215: Doors open, fan off, no sample, DI rinse
061515: Doors closed, fan off, no sample, tested system, power supply failed testing, took parts away to acid bath, while we wait for new power supply.
Plan to mount passive sensor and meteorological sensors this week and install new power supply to active sampler when it arrives.
System is set to come on at RH>91%. Daily maintenance will hopefully mean we can catch the small fog events.
System is now on last year’s program, coming on at RH>91%
Beautiful day down at the LML site! Small fog sample collected today.
Collector deployed today. Fog rolling in at 5 pm. Passive installed without tipping bucket.
No sample, but system came on for 15 minutes. Rinsed collector in place and reused sample jar.
225 ml, clear sample, sunny weather @ 0900 following a heavy fog night, 150608 0900, standard procedure with ion aliquot taken.